OpenStudy (anonymous):

Determine the factors of x2 − 8x − 12. (x − 6)(x + 2) (x + 3)(x − 4) Prime (x + 6)(x − 2)

OpenStudy (anonymous):

Find two numbers that multiply to make your constant and add to make the coefficient of x

OpenStudy (anonymous):

i think is is C

OpenStudy (anonymous):

actually... none of those options seem to work... unless I dont know what im talking about anymore..

hartnn (hartnn):

it does not have integral factors

hartnn (hartnn):

because b^2 -4ac is not a perfect square

OpenStudy (anonymous):

/;gt\

hartnn (hartnn):

you can safely choose prime here....but it has irrational factors

HanAkoSolo (jamierox4ev3r):

so just for this question because of multiple choice, it would be prime, but it does have irrational factors..ha i should have known

HanAkoSolo (jamierox4ev3r):

now wonder i was confused okay :P

hartnn (hartnn):

x = 8 +/- sqrt (108) /2

hartnn (hartnn):

when b^2-4ac is perfect square ----> integer roots b^2-4ac is positive ----->real roots b^2 -4ac is negative ---->imaginary roots --->prime

hartnn (hartnn):

here, if domain of x is integer, only then we can consider it as prime

HanAkoSolo (jamierox4ev3r):

ahh okay then :)

HanAkoSolo (jamierox4ev3r):

very thorough explanation

hartnn (hartnn):

) thanks! i hope harlie got it ...

OpenStudy (debbieg):

Actually, I think the definition of a "prime polynomial" is when it cannot be factored into the product of two polynomials, using integer values. It doesn't mean that it "can't be written as a product" because any trinomial can, with rational, irrational or complex roots. :)

OpenStudy (anonymous):

Nnope i still dont get it

hartnn (hartnn):

which part ?

OpenStudy (anonymous):

the one u said

hartnn (hartnn):

you know what are a,b, c ?

OpenStudy (anonymous):

is the answer A

OpenStudy (debbieg):

Did you check A by doing the FOIL multiplication, to see if you get back the original polynomial?

OpenStudy (debbieg):

You can check each answer by multiplying (one of the shortcomings of this kind of questions being multiple choice, since that doesn't teach you a darn thing about how to factor the polynomial). But as @hartnn points out, the RIGOROUS way to show that it does not factor is to calculate the discriminant. He told you above what that tells you about the roots (and hence, the ability to factor).

OpenStudy (anonymous):

I am not sure its because i only have 3 days to finish what i am doing so i need to know what is the answer

OpenStudy (debbieg):

Do you know how to FOIL multiply?

OpenStudy (debbieg):

(x − 6)(x + 2) = ???

OpenStudy (anonymous):

x2 − 8x − 12.

OpenStudy (debbieg):

How did you get that? show me each of F-O-I-L steps.

OpenStudy (debbieg):

Meaning, show all 4 terms.

OpenStudy (anonymous):

i didn't i just guessed

OpenStudy (debbieg):

Well, therein lies the problem. :) You aren't supposed to guess. Clearly, whatever the class is that you're in, this is material that you are expected to know. So if you don't learn how to do it, you will only get more and more behind. FOIL mutliplication, remember this? 

OpenStudy (debbieg):

So you get (x)(x) + 2(x)-6(x) -6(2)

OpenStudy (anonymous):

is it B

OpenStudy (debbieg):

Did you FOIL b to see if you get back to the original polynomial?

OpenStudy (debbieg):

(x + 3)(x − 4) = ?

OpenStudy (anonymous):

idk

OpenStudy (debbieg):

Did you look at my example above on how to FOIL? (x + 3)(x − 4) = x(x) -4(x) +3(x) -4(3) = ??

OpenStudy (debbieg):

If you don't remember how to FOIL multiply, then you really need to review that before you go any further. You are not prepared to do factoring problems, if you don't understand multiplying first.

OpenStudy (anonymous):

i know how to use the FOIL METHOD

OpenStudy (debbieg):

You can review it here: http://www.coolmath.com/algebra/03-polynomials/06-polynomials-multiplying-foil-01.htm

OpenStudy (debbieg):

Ok, that's good. Then why, when I asked: (x + 3)(x − 4) = ? Did you say "IDK"? Just FOIL it. :)

OpenStudy (anonymous):

x2 − 8x − 12.

OpenStudy (debbieg):

No. Your middle term is wrong.

OpenStudy (debbieg):

(x + 3)(x − 4) = x(x) -4(x) +3(x) -4(3) = ?? -4(x) +3(x)\(\neq\)-8x

OpenStudy (anonymous):

so it the last one

OpenStudy (debbieg):

Did you FOIL (x + 6)(x − 2) to check?

OpenStudy (debbieg):

(x + 6)(x − 2) = ??

OpenStudy (debbieg):

You need to know how to FOIL. IT'S IMPORTANT and if you don't, you will only be more and more lost when you get to the next topic. It's clear from the fact that you are being asked this problem that you are expected to know how to FOIL two binomials, and then how to factor ("un-FOIL") a trinomial. You can get this assignment turned in with every problem correct, but if you don't understand these basic concepts, you'll just be back here even more lost when you have to reduce a rational expression or find its domain, because you don't understand factoring a trinomial.

OpenStudy (anonymous):

The answer prime

OpenStudy (debbieg):

Well, if you've checked the other 3 by correctly FOIL'ing them, and none of them work, then that would be all that's left.